OpenStudy (anonymous):

(-2)/(x+4) = (4)/(x+3)

OpenStudy (phi):

do you know how to cross multiply ?

OpenStudy (anonymous):

I do but lets just go with no

OpenStudy (phi):

a/b = c/d after cross multiplying you get ad = bc what do you get when you cross multiply ?

OpenStudy (anonymous):

4x+16/-2x-6

OpenStudy (phi):

equals not divide 4x+16 = -2x-6 now add 2x to both sides add -16 to both sides

OpenStudy (anonymous):

6x=22

OpenStudy (phi):

-16-6 is -22 (not +22)

OpenStudy (anonymous):

oh ok

OpenStudy (phi):

so it is 6x = -22 to solve for x, divide both sides by 6, and simplify

OpenStudy (anonymous):

x=-3.6